What is the sum of the digits of the first 50 positive odd integers?
Answer by Edwin McCravy(20056)

As you can see from the list below
in the column of first digits there are
5 0's + 5 1's + 5 2's + ... + 5 9's
That's 0+1+2+3+4+5+6+7+8+9 = 45 times 5 or the sum of the
first or tens digits is 225
And as you can see from the list of second digits there is
this:
1+3+5+7+9 = 25
ten times. So that's 25*10 = 250
So the sum of all the digits is 225+250 = 475
01
03
05
07
09
11
13
15
17
19
21
23
25
27
29
31
33
35
37
39
41
43
45
47
49
51
53
55
57
59
61
63
65
67
69
71
73
75
77
79
81
83
85
87
89
91
93
95
97
99
Edwin